OpenStudy (anonymous):

A certain mountain has an elevation of 19,303 feet. In 1917, the glacier on the peak covered 7 acres. By 2005, this glacier had melted to only 1 acre. a) Assume that this glacier melted at a constant rate each year. Find this yearly rate. b) Use your answer for the part (a) to write the linear equation that gives the acreage A of this glacier t years past 1917. The yearly rate of change is ____ acres/year. What is the equation that gives the acreage of the glacier t years after 1917?

OpenStudy (anonymous):

I'm stuck, if I have m=(1-7)/(88/0) -6/88=0.06818. But I am stuck here

hero (hero):

You have 88/0 as part of your calculation which is by default invalid since you can never divide a number by zero.

OpenStudy (anonymous):

I meant 88-0 and it would equal to -6/88

hero (hero):

Okay, that looks right for the slope. I suppose you're trying to figure out the equation now.

OpenStudy (anonymous):

yes

hero (hero):

If we assume 1900 as the the starting point, we can ignore the "19" which leaves (17, 7) and (105, 1) as the two points on an x,y plane. From there you can figure out the equation of the line using the slope -6/88 and one of those points.

hero (hero):

Actually, we better not do that. 1917 should be the starting point.

OpenStudy (anonymous):

Okay so then how would we to get find out how many years at a constant rate did the glacier melt?

hero (hero):

If 1917 is the starting point, then the points are (0,7) and (88,1). The slope would be m = -6/88. Then you can use the point slope formula if you want.

hero (hero):

If 1917 is the starting point, then the points are (0,7) and (88,1). The slope would be m = -6/88. Then you can use the point slope formula if you want.

OpenStudy (anonymous):

the answer I was looking for was the yearly rate and the equation. for example A+?????????

OpenStudy (anonymous):

I meant A=??

OpenStudy (anonymous):

So the equation would be A=(1-7)/(2005-1917) = -6/88